# Parametrized Sharing for Multi-Agent Hybrid DRL for Multiple Multi-Functional RISs-Aided Downlink NOMA Networks

### v1

@@ I Introduction
In the era of the six-generation (6G) wireless communications, the scarcity of spectrum resources has driven researchers to explore more efficient transmission technologies [ref1]. Among them, non-orthogonal multiple access (NOMA) has emerged as a promising solution due to its capability to serve multiple users simultaneously at the same time by frequency resource [ref3]. Compared to the traditional orthogonal multiple access (OMA) mechanism, NOMA exhibits significantly higher spectral efficiency. However, the performance of NOMA is often hindered by challenges such as severe channel fading and inter-user interference, potentially degrading the signal quality and limit its practical deployment [ref4]. To address the issues, reconfigurable intelligent surfaces (RIS) has been proposed as an enabling technology [ref5]. By adjusting the configuration of RIS elements, a virtual line-of-sight (LoS) link can be established to bypass obstacles between the transmitter and receiver and to improve the signal quality for combating the channel fading [ref2]. Although RIS holds promise for next-generation systems, its practical deployment remains constrained by several inherent limitations. Particularly, it can only be operated over a 180-degree half-space coverage area, depending on external power supplies, which confines its independent operation and scalability.
@@
Against the backdrops of RISs, the concept of multi-functional RIS (MF-RIS) has been proposed [ref6], integrating the function of simultaneous transmission and reflection RIS (STAR-RIS) [ref2], providing a 360-degree full-space coverage, realizing a ubiquitous service [ref8], [ref9]. Moreover, energy harvesting (EH) capability at the radio-frequency (RF) is designed in MF-RIS, allowing it to capture wireless energy from incident electromagnetic signals and operating in a self-sustainable manner [ref13]. This design reduces the requirements on the wired power infrastructure or frequent battery replacement, improving system energy efficiency (EE) and deployment flexibility. Additionally, MF-RIS enhances the traditional passive reflection by incorporating active components for signal amplification, improving the weak channel conditions in NOMA networks [ref11, ref12, ref15]. Moreover, there will be increasing needs of deploying multiple MF-RISs for wider coverage requirements. In this work, we investigate a novel architecture of deploying multiple MF-RISs for assisting downlink NOMA networks [ref16]. NOMA shares the same spectrum in multi-MF-RISs-aided networks. On the other hand, MF-RISs contribute to constructing favorable channel conditions for NOMA user groups by mitigating channel fading and interference effects. Moreover, we design based on deep reinforcement learning (DRL) techniques to enable adaptive policy learning under high-dimension and dynamic environments. Unlike conventional DRL handling either discrete or continuous actions separately, a general hybrid DRL framework should be adopted to effectively address complex hybrid continuous-discrete action spaces. The main contributions of this work are summarized as follows:
  • We investigate multi-MF-RISs-aided downlink NOMA networks. We consider power-domain NOMA, where a group of users shares the same frequency resource. MF-RISs are capable of extending the transmission range by reflecting, transmitting, and amplifying signals, while harvesting partial signal energy for operation.
@@
  • We aim at maximizing system EE by deciding power allocation, base station (BS) beamforming and MF-RIS configurations of amplification/phase-shifts/EH ratios and MF-RIS positions. Note that MF-RIS circuit power is also considered. A parametrized sharing in multi-agent hybrid deep reinforcement learning (PMHRL) scheme is designed, whereas hybrid DRL tackles joint continuous-discrete variables respectively by proximal policy optimization (PPO) and deep-Q network (DQN). Parametrized sharing enables information sharing between dual-modules.
  • Results have demonstrated that PMHRL achieves the highest EE compared to other existing benchmarks of conventional DRLs and those without parametrized sharing. The proposed architecture of multi-MF-RISs-aided downlink NOMA achieves the highest EE among the cases without EH, conventional RISs, non-amplified signals and deployments without RISs/MF-RISs.
@@ II System Model and Problem Formulation
In a multi-MF-RISs-assisted downlink NOMA network in Fig. 1, we consider a BS equipped with $N$ transmit antennas with the set of $\mathcal{N}=\{1,2,...,N\}$, serving $J$ users at direction $k$ indexed by the set of $\mathcal{J}_{k}=\{1,2,...,J_{k}\}$. We consider total $K$ direction for NOMA transmission groups, where $\mathcal{K}=\{1,2.....K\}$. We consider $Q$ MF-RISs with its set of $\mathcal{Q}=\{1,2,\ldots,Q\}$. Furthermore, we consider a Cartesian coordinate system with the locations of the BS, MF-RIS, and user being $\mathbf{w}_{b}=[x_{b},y_{b},z_{b}]^{\mathrm{T}}$, $\mathbf{w}_{q}=[x_{q},y_{q},z_{q}]^{\mathrm{T}}$, and $\mathbf{w}_{kj}=[x_{kj},y_{kj},0]^{\mathrm{T}}$, respectively. Note that $\mathrm{T}$ indicates the transpose operation. Due to the limited coverage of MF-RIS, its deployable region is also limited by $\mathcal{W}$, where the following constraint is satisfied: $\mathbf{w}_{q}\in\mathcal{W}=\{[x_{q},y_{q},z_{q}]^{\mathrm{T}}|\mathbf{w}_{\min}\preceq\mathbf{w}_{q}\preceq\mathbf{w}_{\max}\}$ with its deployable areas bounded by $\mathbf{w}_{\min}$ and $\mathbf{w}_{\max}$. Each MF-RIS has $M$ elements indexed by the set of $\mathcal{M}=\{1,2,\ldots,M\}$ with a two-dimensional array with $M=M_{h}\cdot M_{v}$ elements, where $M_{h}$ and $M_{v}$ indicate the respective numbers of elements in horizontal and vertical axes. Each MF-RIS configuration can be defined as $\mathbf{\Theta}_{q}^{k}=\operatorname{diag}\left(\alpha_{q,1}\sqrt{\beta_{q,1}^{k}}e^{j\theta_{q,1}^{k}},\ldots,\alpha_{q,M}\sqrt{\beta_{q,M}^{k}}e^{j\theta_{q,M}^{k}}\right)$, where $\theta_{q,m}^{k}\in[0,2\pi)$ and $\beta_{q,m}^{k}\in[0,\beta^{k}_{\max}]$ denote the phase-shift and amplitude coefficients of MF-RIS at $k$-th direction, respectively. Note that $\beta_{\max}>1$ denotes the signal amplification, whereas $\beta_{\max}\leq 1$ indicates conventional RIS without amplification capability. Each element of the MF-RIS can operate in energy harvesting (EH) mode (H mode) and signal mode (S mode) by adjusting the EH coefficient $\alpha_{q,m}\in\{0,1\}$. Note that $\alpha_{q,m}=1$ implies that MF-RIS operates in S mode, whilst $\alpha_{q,m}=0$ indicates that it functions in only H mode.
@@
We consider the Rician fading channel model between the BS and $q$-th MF-RIS as $\mathbf{H}_{q}=\sqrt{h_{0}d_{q}^{-k_{0}}}\left(\sqrt{\frac{\beta_{0}}{\beta_{0}+1}}\mathbf{H}_{q}^{\text{LoS}}+\sqrt{\frac{1}{\beta_{0}+1}}\mathbf{H}_{q}^{\text{NLoS}}\right)\in\mathbb{C}^{M\times N}$, where $h_{0}$ is the pathloss at the reference distance of 1 meter, $d_{q}=\|\mathbf{w}_{b}-\mathbf{w}_{q}\|^{2}$ is the distance, and $k_{0}$ is the pathloss exponent. $\beta_{0}$ is the Rician factor adjusting the portion of LoS path $\mathbf{H}_{q}^{\text{LoS}}$ and non-LoS (NLoS) component of $\mathbf{H}_{q}^{\text{NLoS}}$. The LoS component [ref19] is expressed as $\mathbf{H}_{q}^{\text{LoS}}=\begin{bmatrix}1,e^{-j\frac{2\pi}{\lambda}d_{R}\sin\bar{\psi}_{r,q}\sin\bar{\theta}_{r,q}},\cdots,e^{-j\frac{2\pi}{\lambda}(M_{z}\!-\!1)d_{R}\sin\bar{\psi}_{r,q}\sin\bar{\theta}_{r,q}}\end{bmatrix}^{\mathrm{T}}\otimes\begin{bmatrix}1,e^{-j\frac{2\pi}{\lambda}d_{R}\sin\bar{\psi}_{r,q}\cos\bar{\theta}_{r,q}},\cdots,e^{-j\frac{2\pi}{\lambda}(M_{y}\!-\!1)d_{R}\sin\bar{\psi}_{r,q}\cos\bar{\theta}_{r,q}}\end{bmatrix}^{\mathrm{T}}\!\otimes\!\begin{bmatrix}1,e^{-j\frac{2\pi}{\lambda}d_{B}\sin\varphi_{t}\cos\vartheta_{t}},\cdots,e^{-j\frac{2\pi}{\lambda}(N\!-\!1)d_{B}\sin\varphi_{t}\cos\vartheta_{t}}\end{bmatrix}^{\mathrm{T}}$, where $\otimes$ denotes the Kronecker product and $T$ is transpose operation. $\lambda$ indicates the wavelength of the operating frequency. Notations of $d_{R}$ and $d_{B}$ denote the element spacing of MF-RIS and antenna separation of BS, respectively. Notations of $\bar{\psi}_{r,q}$, $\bar{\theta}_{r,q}$, $\varphi_{t}$, and $\vartheta_{t}$ represent the azimuth and elevation angles of-arrivals of MF-RIS $q$, and those of angle-of-departures of BS, respectively. Note that $\mathbf{H}^{\text{NLoS}}_{q}$ follows the Rayleigh fading. The direct link from BS and reflected link from the MF-RIS $q$ to user $j$ at direction $k$ are denoted by $\mathbf{h}_{kj}\in\mathbb{C}^{N\times 1}$ and $\mathbf{r}_{q,kj}\in\mathbb{C}^{M\times 1}$, respectively, associated with their distances of $d_{kj}$ and $d_{q,kj}$. While, both parameters follow $\mathbf{H}_{q}$ but in a vector form, where the LoS components are $\mathbf{h}^{\text{LoS}}_{kj}\!=\![1,e^{-j\frac{2\pi}{\lambda}d_{B}\sin\varphi_{t}\sin\vartheta_{t}},\cdots,e^{-j\frac{2\pi}{\lambda}(N\!-\!1)d_{B}\sin\varphi_{t}\sin\vartheta_{t}}]^{\mathrm{T}}$ and $\mathbf{r}_{q,kj}^{\text{LoS}}=[1,e^{-j\!\frac{2\pi}{\lambda}\!d_{R}\sin\!\varphi_{t,q}\sin\!\vartheta_{t,q}},\cdots,e^{-j\!\frac{2\pi}{\lambda}\!(M\!-\!1)d_{R}\sin\!\varphi_{t,q}\sin\!\vartheta_{t,q}}]^{\mathrm{T}}$. The NLoS parts $\mathbf{h}_{kj}^{\text{NLoS}}$ and $\mathbf{r}_{q,kj}^{\text{NLoS}}$ are both characterized by Rayleigh fading. Accordingly, the channel between the MF-RIS $q$ and and user $j$ at direction $k$ is $\mathbf{g}_{q,kj}=\mathbf{r}_{q,kj}^{H}\mathbf{\Theta}_{q}^{k}\mathbf{H}_{q}$ where $H$ indicates Hermitian operation. The total combined channel of BS-user $j$ at direction $k$ assisted by $Q$ MF-RISs is defined as ${\mathbf{g}}_{kj}=\mathbf{h}_{kj}+\sum_{q\in\mathcal{Q}}\mathbf{g}_{q,kj}$.
@@
where $\mathbf{f}_{k}$ represents the transmit beamforming vector of the BS for direction $k$. Moreover, $p_{kj}$ denotes the power allocation factor for user $j$ at direction $k$ where $\sum_{j\in\mathcal{J}_{k}}p_{kj}=1$. $\mathbf{n}_{q}\sim\mathcal{CN}(0,\sigma_{s}^{2}\mathbf{I}_{M})$ denotes the amplification noise from MF-RISs with element noise power $\sigma_{s}^{2}$. Notation of $n_{kj}$ is noise power of user $j$ at direction $k$ with its power of $\sigma_{u}^{2}$. Moreover, NOMA user signals are transmitted simultaneously at the same frequency, leading to mutual interference. To decode the intended signals, users employ successively interference cancellation (SIC) [ref14]. Assume that the users $j$ and $l$ in direction $k$ are ranked in an ascending order according to the equivalent combined channel gains, associated with the conditions of
@@
where $k\in\mathcal{K}$, $j\in\mathcal{J}_{k}$ denote users at direction $k$, and $l\in\mathcal{L}_{k}=\{j,j+1,\cdots,J_{k}\}$. Notation $I_{kj}=\sum_{q\in\mathcal{Q}}\sigma_{s}^{2}\|\mathbf{r}_{q,kj}^{H}\mathbf{\Theta}_{q}^{k}\|^{2}$ indicates the residual interference. The signal-to-interference-plus-noise ratio (SINR) is given by
$\gamma_{kj}=\frac{|\mathbf{g}_{kl}\mathbf{f}_{k}|^{2}p_{kj}}{\sum_{l\in\mathcal{L}_{k}}|\mathbf{g}_{kl}\mathbf{f}_{k}|^{2}p_{kl}+I_{\text{IG},k}+I_{\text{MR}}+\sigma_{u}^{2}},$ (3)
where $I_{\text{IG},k}=\sum_{\bar{k}\in\mathcal{K}\setminus\{k\}}\sum_{i\in\mathcal{J}_{\bar{k}}}\|\mathbf{g}_{kj}\mathbf{f}_{\bar{k}}\|^{2}p_{\bar{k}i}$ indicates the inter-group interference, and $I_{\text{MR}}=\sum_{q\in\mathcal{Q}}\sigma_{s}^{2}\|\mathbf{r}_{q,kj}^{H}\mathbf{\Theta}_{q}^{k}\|^{2}$ denotes the noise induced from multi-MF-RISs. Therefore, the achievable rate for user $j$ at direction $k$ can be expressed as $R_{kj}=\log_{2}(1+\gamma_{kj})$.
@@
Here, we define the EH coefficient matrix for the $m$-th element of the $q$-th MF-RIS as $\mathbf{T}_{q,m}=\operatorname{diag}\left(\left[{0,\dots,0},1-\alpha_{q,m},{0,\dots,0}\right]\right)$. Therefore, the RF power recieved by the $m$-th element of the $q$-th MF-RIS is given by $P_{q,m}^{\text{RF}}=\mathbb{E}\left(\left\|\mathbf{T}_{q,m}\left(\mathbf{H}_{q}\sum_{k\in\mathcal{K}}\mathbf{f}_{k}+\mathbf{n}_{q,m}\right)\right\|^{2}\right)$, where $\mathbf{n}_{q,m}$ is the amplified noise introduced by the MF-RIS. To capture RF energy conversion efficiency for different input power, a non-linear harvesting model is adopted. Accordingly, the total power of the $m$-th element of the $q$-th MF-RIS is expressed as $P_{q,m}^{\text{A}}=\frac{\Upsilon_{q,m}-Z\Omega}{1-\Omega}$, where $\Upsilon_{q,m}=\frac{Z}{1+e^{-p\left(P_{q,m}^{\text{RF}}-k\right)}}$ is a logistic function with respect to the received RF power $P^{\text{RF}}_{q,m}$, and $Z>0$ is a constant determining the maximum harvested power. Constant $\Omega=\frac{1}{1+e^{\varpi_{1}\varpi_{2}}}$ ensures a zero-input/zero-output response in H mode with constants $\varpi_{1}>0$ and $\varpi_{2}>0$ capturing the effects of circuit sensitivity and current leakage. To achieve the self-sustainability, the total consumed power of MF-RISs should be lower than the harvested power. Moreover, the power for controlling MF-RIS mainly comes from the total number of PIN diodes required [ref17-1]. The quantization levels assigned for EH ratio, amplitude and phase shifts are $L_{\alpha}$, $L_{\beta}$, and $L_{\theta}$, respectively, where the total number of PIN diodes per MF-RIS is $\log_{2}L_{\alpha}+K\log_{2}L_{\beta}+K\log_{2}L_{\theta}$. We have the following self-sustainability constraint per MF-RIS, i.e., $P_{q}^{\text{con}}\leq\sum_{m\in\mathcal{M}}P_{q,m}^{\text{A}}$, where $P_{q}^{\text{con}}=\lceil\log_{2}L_{\alpha}+K\log_{2}L_{\beta}+K\log_{2}L_{\theta}\rceil\cdot MP_{\text{PIN}}+P_{C}+\xi\cdot P_{l,O}$. Here, $P_{C}$ denotes the power consumption of RF-to-DC power conversion, and $P_{\text{PIN}}$ is power consumption per PIN diode. Notation $\xi$ indicates the inverse of amplifier efficiency. The output power of MF-RIS $q$ is obtained as $P_{O,q}=\sum_{k\in\mathcal{K}}\left(\sum_{k^{\prime}\in\mathcal{K}}\|\boldsymbol{\Theta}_{q}^{k}\mathbf{H}_{q}\mathbf{f}_{k^{\prime}}\|^{2}+\sigma_{s}^{2}\|\boldsymbol{\Theta}_{q}^{k}\|_{F}^{2}\right)$, where $\|\cdot\|_{F}$ is Forbenius norm.

### v2

@@ I Introduction
In the era of the six-generation (6G) wireless communications, the scarcity of spectrum resources has driven researchers to explore more efficient transmission technologies [10]. Among them, non-orthogonal multiple access (NOMA) has emerged as a promising solution due to its capability to serve multiple users simultaneously at the same time by frequency resource [6]. Compared to the traditional orthogonal multiple access (OMA) mechanism, NOMA exhibits significantly higher spectral efficiency. However, the performance of NOMA is often hindered by challenges such as severe channel fading and inter-user interference, potentially degrading the signal quality and limit its practical deployment [20]. To address the issues, reconfigurable intelligent surfaces (RIS) has been proposed as an enabling technology [12]. By adjusting the configuration of RIS elements, a virtual line-of-sight (LoS) link can be established to bypass obstacles between the transmitter and receiver and to improve the signal quality for combating the channel fading. Although RIS holds promise for next-generation systems, its practical deployment remains constrained by several inherent limitations. Particularly, it can only be operated over a 180-degree half-space coverage area, depending on external power supplies, which confines its independent operation and scalability.
@@
Against the backdrops of RISs, the concept of multi-functional RIS (MF-RIS) has been proposed [18, 13], integrating the function of simultaneous transmission and reflection RIS (STAR-RIS), providing a 360-degree full-space coverage, realizing a ubiquitous service [3, 15]. The work of [18] has established the fundamental system modeling and corresponding performance optimization of the MF-RIS deployment. Moreover, energy harvesting (EH) capability at the radio-frequency (RF) is designed in MF-RIS, allowing it to capture wireless energy from incident electromagnetic signals and operating in a self-sustainable manner [2, 5]. This design reduces the requirements on the wired power infrastructure or frequent battery replacement, improving system energy efficiency (EE) and deployment flexibility [14]. The authors in [2] also conduct sum-rate maximization with robust and secure transmissions by MF-RIS. The work of [17] has analyzed the feasibility and its performance when leveraging MF-RIS with self-sustainable capability. Additionally, MF-RIS enhances the traditional passive reflection by incorporating active components for signal amplification, improving the weak channel conditions in NOMA networks [4, 22, 16]. Moreover, there will be increasing needs of deploying multiple MF-RISs for wider coverage requirements. Extending from prior work of [18] with signal MF-RIS and rate optimization, we investigate a novel architecture of deploying multiple MF-RISs for assisting downlink NOMA networks [7], associated with the EE optimization. NOMA shares the same spectrum in multi-MF-RISs-aided networks. On the other hand, MF-RISs contribute to constructing favorable channel conditions for NOMA user groups by mitigating channel fading and interference effects. Moreover, we design based on deep reinforcement learning (DRL) [11] techniques to enable adaptive policy learning under high-dimension and dynamic environments. Unlike conventional DRL handling either discrete or continuous actions separately, a general hybrid DRL framework should be adopted to effectively address complex hybrid continuous-discrete action spaces.
@@
The main contributions of this work are summarized as follows: (1) We investigate multi-MF-RISs-aided downlink NOMA networks. We consider power-domain NOMA, where a group of users shares the same frequency resource. MF-RISs are capable of extending the transmission range by reflecting, transmitting, and amplifying signals, while harvesting partial signal energy for operation. (2) We aim at maximizing system EE by deciding power allocation, base station (BS) beamforming and MF-RIS configurations of amplification/phase-shifts/EH ratios and MF-RIS positions. Note that MF-RIS circuit power is also considered. A parametrized sharing in multi-agent hybrid deep reinforcement learning (PMHRL) scheme is designed, whereas hybrid DRL tackles joint continuous-discrete variables respectively by proximal policy optimization (PPO) and deep-Q network (DQN). Parametrized sharing enables information sharing between dual-modules. (3) Results have demonstrated that PMHRL achieves the highest EE compared to other existing benchmarks of conventional DRLs and those without parametrized sharing. The proposed architecture of multi-MF-RISs-aided downlink NOMA achieves the highest EE among the cases without EH, conventional RISs, non-amplified signals and deployments without RISs/MF-RISs.
@@ II System Model and Problem Formulation
In a multi-MF-RISs-assisted downlink NOMA network in Fig. 1, we consider a BS equipped with $N$ transmit antennas with the set of $\mathcal{N}=\{1,2,...,N\}$, serving $J$ users at direction $k$ indexed by the set of $\mathcal{J}_{k}=\{1,2,...,J_{k}\}$. We consider total $K$ direction for NOMA transmission groups, where $\mathcal{K}=\{1,2.....K\}$. We consider $Q$ MF-RISs with its set of $\mathcal{Q}=\{1,2,\ldots,Q\}$. Furthermore, we consider a Cartesian coordinate system with the locations of the BS, MF-RIS, and user being $\mathbf{w}_{b}=[x_{b},y_{b},z_{b}]^{\mathrm{T}}$, $\mathbf{w}_{q}=[x_{q},y_{q},z_{q}]^{\mathrm{T}}$, and $\mathbf{w}_{kj}=[x_{kj},y_{kj},0]^{\mathrm{T}}$, respectively. Note that $\mathrm{T}$ indicates the transpose operation. Due to the limited coverage of MF-RIS, its deployable region is also limited by $\mathcal{W}$, where the following constraint is satisfied: $\mathbf{w}_{q}\in\mathcal{W}=\{[x_{q},y_{q},z_{q}]^{\mathrm{T}}|\mathbf{w}_{\min}\preceq\mathbf{w}_{q}\preceq\mathbf{w}_{\max}\}$ with its deployable areas bounded by $\mathbf{w}_{\min}$ and $\mathbf{w}_{\max}$. Each MF-RIS has $M$ elements indexed by the set of $\mathcal{M}=\{1,2,\ldots,M\}$ with a two-dimensional array with $M=M_{h}\cdot M_{v}$ elements, where $M_{h}$ and $M_{v}$ indicate the respective numbers of elements in horizontal and vertical axes. Each MF-RIS configuration can be defined as $\mathbf{\Theta}_{q}^{k}=\operatorname{diag}\left(\alpha_{q,1}\sqrt{\beta_{q,1}^{k}}e^{j\theta_{q,1}^{k}},\ldots,\alpha_{q,M}\sqrt{\beta_{q,M}^{k}}e^{j\theta_{q,M}^{k}}\right)$, where $\theta_{q,m}^{k}\in[0,2\pi)$ and $\beta_{q,m}^{k}\in[0,\beta^{k}_{\max}]$ denote the phase-shift and amplitude coefficients of MF-RIS at $k$-th direction, respectively. Note that $\beta_{\max}>1$ denotes the signal amplification, whereas $\beta_{\max}\leq 1$ indicates conventional RIS without amplification capability. Each element of the MF-RIS can operate in harvesting mode (H mode) and signal mode (S mode) by adjusting the EH coefficient $\alpha_{q,m}\in\{0,1\}$. Note that $\alpha_{q,m}=1$ implies that MF-RIS operates in S mode, whilst $\alpha_{q,m}=0$ indicates that it functions in only H mode.
@@
We consider the Rician fading channel model between the BS and $q$-th MF-RIS as $\mathbf{H}_{q}=\sqrt{h_{0}d_{q}^{-k_{0}}}\left(\sqrt{\frac{\beta_{0}}{\beta_{0}+1}}\mathbf{H}_{q}^{\text{LoS}}+\sqrt{\frac{1}{\beta_{0}+1}}\mathbf{H}_{q}^{\text{NLoS}}\right)\in\mathbb{C}^{M\times N}$, where $h_{0}$ is the pathloss at the reference distance of 1 meter, $d_{q}=\|\mathbf{w}_{b}-\mathbf{w}_{q}\|^{2}$ is the distance, and $k_{0}$ is the pathloss exponent. $\beta_{0}$ is the Rician factor adjusting the portion of LoS path $\mathbf{H}_{q}^{\text{LoS}}$ and non-LoS (NLoS) component of $\mathbf{H}_{q}^{\text{NLoS}}$. The LoS component [18] is expressed as $\mathbf{H}_{q}^{\text{LoS}}=\begin{bmatrix}1,e^{-j\frac{2\pi}{\lambda}d_{R}\sin\bar{\psi}_{r,q}\sin\bar{\theta}_{r,q}},\cdots,e^{-j\frac{2\pi}{\lambda}(M_{z}\!-\!1)d_{R}\sin\bar{\psi}_{r,q}\sin\bar{\theta}_{r,q}}\end{bmatrix}^{\mathrm{T}}\otimes\begin{bmatrix}1,e^{-j\frac{2\pi}{\lambda}d_{R}\sin\bar{\psi}_{r,q}\cos\bar{\theta}_{r,q}},\cdots,e^{-j\frac{2\pi}{\lambda}(M_{y}\!-\!1)d_{R}\sin\bar{\psi}_{r,q}\cos\bar{\theta}_{r,q}}\end{bmatrix}^{\mathrm{T}}\!\otimes\!\begin{bmatrix}1,e^{-j\frac{2\pi}{\lambda}d_{B}\sin\varphi_{t}\cos\vartheta_{t}},\cdots,e^{-j\frac{2\pi}{\lambda}(N\!-\!1)d_{B}\sin\varphi_{t}\cos\vartheta_{t}}\end{bmatrix}^{\mathrm{T}}$, where $\otimes$ denotes the Kronecker product and $T$ is transpose operation. $\lambda$ indicates the wavelength of the operating frequency. Notations of $d_{R}$ and $d_{B}$ denote the element spacing of MF-RIS and antenna separation of BS, respectively. Notations of $\bar{\psi}_{r,q}$, $\bar{\theta}_{r,q}$, $\varphi_{t}$, and $\vartheta_{t}$ represent the azimuth and elevation angles of-arrivals of MF-RIS $q$, and those of angle-of-departures of BS, respectively. Note that $\mathbf{H}^{\text{NLoS}}_{q}$ follows the Rayleigh fading. The direct link from BS and reflected link from the MF-RIS $q$ to user $j$ at direction $k$ are denoted by $\mathbf{h}_{kj}\in\mathbb{C}^{N\times 1}$ and $\mathbf{r}_{q,kj}\in\mathbb{C}^{M\times 1}$, respectively, associated with their distances of $d_{kj}$ and $d_{q,kj}$. While, both parameters follow $\mathbf{H}_{q}$ but in a vector form, where the LoS components are $\mathbf{h}^{\text{LoS}}_{kj}\!=\![1,e^{-j\frac{2\pi}{\lambda}d_{B}\sin\varphi_{t}\sin\vartheta_{t}},\cdots,e^{-j\frac{2\pi}{\lambda}(N\!-\!1)d_{B}\sin\varphi_{t}\sin\vartheta_{t}}]^{\mathrm{T}}$ and $\mathbf{r}_{q,kj}^{\text{LoS}}=[1,e^{-j\!\frac{2\pi}{\lambda}\!d_{R}\sin\!\varphi_{t,q}\sin\!\vartheta_{t,q}},\cdots,e^{-j\!\frac{2\pi}{\lambda}\!(M\!-\!1)d_{R}\sin\!\varphi_{t,q}\sin\!\vartheta_{t,q}}]^{\mathrm{T}}$. The NLoS parts $\mathbf{h}_{kj}^{\text{NLoS}}$ and $\mathbf{r}_{q,kj}^{\text{NLoS}}$ are both characterized by Rayleigh fading. Accordingly, the channel between the MF-RIS $q$ and and user $j$ at direction $k$ is $\mathbf{g}_{q,kj}=\mathbf{r}_{q,kj}^{H}\mathbf{\Theta}_{q}^{k}\mathbf{H}_{q}$ where $H$ indicates Hermitian operation. The total combined channel of BS-user $j$ at direction $k$ assisted by $Q$ MF-RISs is defined as ${\mathbf{g}}_{kj}=\mathbf{h}_{kj}+\sum_{q\in\mathcal{Q}}\mathbf{g}_{q,kj}$.
@@
where $\mathbf{f}_{k}$ represents the transmit beamforming vector of the BS for direction $k$. Moreover, $p_{kj}$ denotes the power allocation factor for user $j$ at direction $k$ where $\sum_{j\in\mathcal{J}_{k}}p_{kj}=1$. $\mathbf{n}_{q}\sim\mathcal{CN}(0,\sigma_{s}^{2}\mathbf{I}_{M})$ denotes the amplification noise from MF-RISs with element noise power $\sigma_{s}^{2}$. Notation of $n_{kj}$ is noise power of user $j$ at direction $k$ with its power of $\sigma_{u}^{2}$. Moreover, NOMA user signals are transmitted simultaneously at the same frequency, leading to mutual interference. To decode the intended signals, users employ successively interference cancellation (SIC) [8]. Assume that the users $j$ and $l$ in direction $k$ are ranked in an ascending order according to the equivalent combined channel gains, i.e., $|{\mathbf{g}}_{k1}|^{2}\geq...\geq|{\mathbf{g}}_{kj}|^{2}\geq|{\mathbf{g}}_{kJ}|^{2}$, associated with the SIC conditions of
@@
where $k\in\mathcal{K}$, $j\in\mathcal{J}_{k}$ denote users at direction $k$, and $l\in\mathcal{L}_{k}=\{j,j+1,\cdots,J_{k}\}$. Notation $I_{kj}=\sum_{q\in\mathcal{Q}}\sigma_{s}^{2}\|\mathbf{r}_{q,kj}^{H}\mathbf{\Theta}_{q}^{k}\|^{2}$ indicates the residual interference. Note that inter-/intra-group interferences are considered constants and is negligible in SIC conditions in (2). However, such effects should be taken into account in calculating signal-to-interference-plus-noise ratio (SINR) given by $\gamma_{kj}=\frac{|\mathbf{g}_{kl}\mathbf{f}_{k}|^{2}p_{kj}}{\sum_{l\in\mathcal{L}_{k}}|\mathbf{g}_{kl}\mathbf{f}_{k}|^{2}p_{kl}+I_{\text{IG},k}+I_{kj}+\sigma_{u}^{2}}$, where $I_{\text{IG},k}=\sum_{\bar{k}\in\mathcal{K}\setminus\{k\}}\sum_{i\in\mathcal{J}_{\bar{k}}}\|\mathbf{g}_{kj}\mathbf{f}_{\bar{k}}\|^{2}p_{\bar{k}i}$ indicates the inter-group interference. Note that the first term of the denominator of SINR means the intra-group interference after SIC. Therefore, the achievable rate for user $j$ at direction $k$ can be expressed as $R_{kj}=\log_{2}(1+\gamma_{kj})$.
@@
Here, we define the EH coefficient matrix for the $m$-th element of the $q$-th MF-RIS as $\mathbf{T}_{q,m}=\operatorname{diag}\left(\left[{0,\dots,0},1-\alpha_{q,m},{0,\dots,0}\right]\right)$. Therefore, the RF power recieved by the $m$-th element of the $q$-th MF-RIS is given by $P_{q,m}^{\text{RF}}=\mathbb{E}\left(\left\|\mathbf{T}_{q,m}\left(\mathbf{H}_{q}\sum_{k\in\mathcal{K}}\mathbf{f}_{k}+\mathbf{n}_{q,m}\right)\right\|^{2}\right)$, where $\mathbf{n}_{q,m}$ is the amplified noise introduced by the MF-RIS. To capture RF energy conversion efficiency for different input power, a non-linear harvesting model is adopted. Accordingly, the total power of the $m$-th element of the $q$-th MF-RIS is expressed as $P_{q,m}^{\text{A}}=\frac{\Upsilon_{q,m}-Z\Omega}{1-\Omega}$, where $\Upsilon_{q,m}=\frac{Z}{1+e^{-p\left(P_{q,m}^{\text{RF}}-k\right)}}$ is a logistic function with respect to the received RF power $P^{\text{RF}}_{q,m}$, and $Z>0$ is a constant determining the maximum harvested power. Constant $\Omega=\frac{1}{1+e^{\varpi_{1}\varpi_{2}}}$ ensures a zero-input/zero-output response in H mode with constants $\varpi_{1}>0$ and $\varpi_{2}>0$ capturing the effects of circuit sensitivity and current leakage. Moreover, the power for controlling MF-RIS mainly comes from the total number of PIN diodes required [19]. The quantization levels assigned for EH ratio, amplitude and phase shifts are $L_{\alpha}$, $L_{\beta}$, and $L_{\theta}$, respectively, where the total number of PIN diodes per MF-RIS is $\log_{2}L_{\alpha}+K\log_{2}L_{\beta}+K\log_{2}L_{\theta}$. We have the following self-sustainability constraint per MF-RIS, i.e., $P_{q}^{\text{con}}\leq\sum_{m\in\mathcal{M}}P_{q,m}^{\text{A}}$, where $P_{q}^{\text{con}}=\lceil\log_{2}L_{\alpha}+K\log_{2}L_{\beta}+K\log_{2}L_{\theta}\rceil\cdot MP_{\text{PIN}}+P_{C}+\xi\cdot P_{l,O}$. Here, $P_{C}$ denotes the power consumption of RF-to-DC power conversion, and $P_{\text{PIN}}$ is power consumption per PIN diode. Notation $\xi$ indicates the inverse of amplifier efficiency. The output power of MF-RIS $q$ is obtained as $P_{O,q}=\sum_{k\in\mathcal{K}}\left(\sum_{k^{\prime}\in\mathcal{K}}\|\boldsymbol{\Theta}_{q}^{k}\mathbf{H}_{q}\mathbf{f}_{k^{\prime}}\|^{2}+\sigma_{s}^{2}\|\boldsymbol{\Theta}_{q}^{k}\|_{F}^{2}\right)$, where $\|\cdot\|_{F}$ is Forbenius norm.